# COSMO-Bench: A Benchmark for Collaborative SLAM Optimization

### v1

@@ Abstract
Recent years have seen a focus on research into distributed optimization algorithms for multi-robot Collaborative Simultaneous Localization and Mapping (C-SLAM). Research in this domain, however, is made difficult by a lack of standard benchmark datasets. Such datasets have been used to great effect in the field of single-robot SLAM, and researchers focused on multi-robot problems would benefit greatly from dedicated benchmark datasets. To address this gap, we design and release the Collaborative Open-Source Multi-robot Optimization Benchmark (COSMO-Bench) – a suite of 24 datasets derived from a state-of-the-art C-SLAM front-end and real-world LiDAR data. Data DOI:  10.1184/R1/29652158. This work was partially supported by NASA award 80NSSC24CA020 and the NSF Graduate Research Fellowship Program. The authors are with the Robotics Institute, Carnegie Mellon University, Pittsburgh, PA, USA. {danmcgann, potokar, kaess}@cmu.edu
@@ I Introduction
In this work we develop the Collaborative Open-Source Multi-robot Optimization Benchmark (COSMO-Bench) [^1] Hosted in perpetuity at  doi.org/10.1184/R1/29652158 and accessible with a user-friendly interface at cosmobench.com. – a suite of 24 benchmark datasets for distributed C-SLAM back-end evaluation (Fig. 1). We work to make COSMO-Bench as realistic as possible by deriving the datasets from real-world sensor data, a state-of-the-art LiDAR-based C-SLAM front-end, and a communication model developed from actual multi-robot communication data.
@@ IV-C Communication Model
Figure 3: Summary of communication data recorded by Lajoie et al. [38], including per-channel throughput (left) and probability of inter-robot connectivity (right) against inter-robot distance. In the connectivity plots, we additionally plot $99\%$ confidence intervals, though these are overly optimistic as the data is correlated.
From these figures we can draw some conclusions about the communication network. Firstly, we can observe that bandwidth is uncorrelated with inter-robot distance; rather, when robots are able to connect, they appear to utilize all the available bandwidth. However, we can see that the bandwidth is not equally shared; we hypothesize that this is a side effect of running multiple instances of network profiles as each attempts to flood the network. When used for actual inter-robot communication, we predict that better network sharing is possible. Second, while bandwidth is not correlated with inter-robot distance, connectivity is correlated. Notably, connectivity between robots is probable at close range and unlikely at larger inter-robot distances.
@@
Using the analysis of this data along with discussions with subject matter experts, we defined a generic communication model that can be used to simulate inter-robot communication.
We assume that all robots engage in a single peer-to-peer communication at a time (to minimize interference) on a shared channel (e.g. Wi-Fi, UWB, acoustic) that has a maximum bandwidth $\mathcal{B}$. Simultaneous communications will cause interference and require bandwidth sharing, which we assume to be allocated equally. However, this interference will be spatially localized due to limited communication ranges. Specifically for a communication between robots $a$ and $b$, there will be interference if there exists an actively communicating robot within a distance $d_{intf}$ meters of either $a$ or $b$. We define the total number of communications occurring around $a$ and $b$ at timestep $t$ as $\mathcal{I}_{ab}(t,\Theta^{*})$, which uses the reference solution $\Theta^{*}$ to compute inter-robot distances. From this we can define the available bandwidth between robots $a$ and $b$ as $\mathcal{B}_{ab}=\mathcal{B}/\mathcal{I}_{ab}(t,\Theta^{*})$.
@@
$\mathcal{T}_{ab}(t)=\left(\mathcal{B}/\mathcal{I}_{ab}(t,\Theta^{*})\right)\cdot\mathcal{C}_{ab}(t,\Theta^{*})\cdot\delta$ (2)
@@
Based on the Wi-Fi model, we also define a model based on “Professional” radio equipment. Such hardware is expected to have increased range but lower bandwidth due to the increased power requirements. The remaining hyperparameters are scaled from those of the Wi-Fi model to maintain a similar structure. Both the Wi-Fi and Pro-Radio models are summarized in Fig. 4.

### v2

@@ Abstract
Recent years have seen a focus on research into distributed optimization algorithms for multi-robot Collaborative Simultaneous Localization and Mapping (C-SLAM). Research in this domain, however, is made difficult by a lack of standard benchmark datasets. Such datasets have been used to great effect in the field of single-robot SLAM, and researchers focused on multi-robot problems would benefit greatly from dedicated benchmark datasets. To address this gap, we design and release the Collaborative Open-Source Multi-robot Optimization Benchmark (COSMO-Bench) – a suite of 24 datasets derived from a baseline C-SLAM front-end and real-world LiDAR data. Data DOI:  10.1184/R1/29652158. This work was partially supported by NASA award 80NSSC24CA020 and the NSF Graduate Research Fellowship Program. The authors are with the Robotics Institute, Carnegie Mellon University, Pittsburgh, PA, USA. {danmcgann, potokar, kaess}@cmu.edu
@@ I Introduction
In this work we develop the Collaborative Open-Source Multi-robot Optimization Benchmark (COSMO-Bench) [^1] Hosted in perpetuity at  doi.org/10.1184/R1/29652158 and accessible with a user-friendly interface at cosmobench.com. – a suite of 24 benchmark datasets for distributed C-SLAM back-end evaluation (Fig. 1). We work to make COSMO-Bench as realistic as possible by deriving the datasets from real-world sensor data, a baseline LiDAR-based C-SLAM front-end, and a communication model developed from actual multi-robot communication data.
@@ IV-C Communication Model
Figure 3: Summary of Wi-Fi communication data recorded by Lajoie et al. [38], including per-channel throughput (left) and probability of inter-robot connectivity (right) against inter-robot distance. In the connectivity plots, we additionally plot $99\%$ confidence intervals, though these are overly optimistic as the data is correlated.
From these figures we can draw some conclusions about the communication network. Firstly, we can observe that bandwidth is uncorrelated with inter-robot distance; rather, when robots are able to connect, they appear to utilize all the available bandwidth. However, we can see that the bandwidth is not equally shared; we hypothesize that this is a side effect of running multiple instances of network profilers as each attempts to flood the network. When used for actual inter-robot communication, we predict that better network sharing is possible. Second, while bandwidth is not correlated with inter-robot distance, connectivity is correlated. Notably, connectivity between robots is probable at close range and unlikely at larger inter-robot distances.
@@
Using the analysis of this data, we defined a generic communication model that can be used to simulate inter-robot communication.
We assume that all robots engage in a single peer-to-peer communication at a time (to minimize interference) on a shared channel (e.g. Wi-Fi, UWB, acoustic) that has a maximum bandwidth $\mathcal{B}$. Simultaneous communications will cause interference and require bandwidth sharing, which we assume to be allocated equally. However, this interference will be spatially localized due to limited communication ranges. Specifically for a communication between robots $a$ and $b$, there will be interference if there exists an actively communicating robot within a distance $d_{intf}$ meters of either $a$ or $b$. We define the total number of communications occurring around $a$ and $b$ at timestep $t$ as $\mathcal{I}_{ab}(t,\Theta^{*})$, which uses the reference solution $\Theta^{*}$ to compute inter-robot distances. From this we can define the available bandwidth between robots $a$ and $b$ as $\mathcal{B}/\mathcal{I}_{ab}(t,\Theta^{*})$.
@@
$\mathcal{T}_{ab}(t)=\underbracket{\left(\mathcal{B}/\mathcal{I}_{ab}(t,\Theta^{*})\right)}_{\mathrm{Bandwidth}}\cdot\underbracket{\mathcal{C}_{ab}(t,\Theta^{*})}_{\mathrm{Connectivity}}\cdot~\delta$ (2)
@@
Based on the Wi-Fi model, we also define a model based on “Professional” radio equipment (Pro-Radio). Such hardware is expected to have increased range but lower bandwidth due to the increased power requirements. The remaining hyperparameters are scaled from those of the Wi-Fi model to maintain a similar structure. Both the Wi-Fi and Pro-Radio models are summarized in Fig. 4.